OpenStudy (anonymous):

If 4/5 of a number is 20, what is 1/2 of the number?

OpenStudy (anonymous):

So you have 4/5x = 20 and yx=0.5

OpenStudy (shamil98):

4/5 x = 20 solve for x. then divide it by 2

OpenStudy (anonymous):

Oops I mean x/2=y

OpenStudy (anonymous):

Do you understand the problem this way or do yo need more help with understanding it?

OpenStudy (anonymous):

The answer is 12.5 as 20 is the 4/5 of 25 in half.

OpenStudy (anonymous):

Yes girl but the purpose of this website is not to hand out answers, but to help people understand why the answer is the answer.